Dry hydrogen by pellet of mercury in a uniform capillary tube closed at one end. If the length of the column of hydrogen at 27 is 1.0 m, at what temperature will the length be 1.20 m?
1
Expert's answer
2020-06-01T14:28:38-0400

Assuming the pressure is constant, write the equality for the two stater of the gas:

"\frac{V_1}{T_1}=\frac{V_2}{T_2},\\\space\\
T_2=T_1\frac{V_2}{V_1}=T_1\frac{Ah_2}{Ah_1}=T_1\frac{h_2}{h_1}=\\\space\\=(273+27)\frac{1.2A}{1A}=360\text{ K},"

or 87°C